OpenStudy (anonymous):

Which is the formula to calculate area for a circle with diameter sixty-eight miles? two pi times thirty-four pi times thirty-four squared two pi times sixty-eight pi times sixty-eight squared

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ARMYRANGERS

OpenStudy (paki):

have you any idea here...?

OpenStudy (paki):

use the formula... A=πd^2/4

OpenStudy (anonymous):

how?

OpenStudy (paki):

what is diameter here... ?

OpenStudy (paki):

@Ail.A please respond

OpenStudy (paki):

ok see here... A=πd^2/4 A=π(68)^2/4

OpenStudy (anonymous):

ok...

OpenStudy (paki):

is this clear now...? @Ail.A

OpenStudy (anonymous):

what does ^ mean

OpenStudy (paki):

mean you got the solution... ?

OpenStudy (anonymous):

no

OpenStudy (tkhunny):

Horrible question. There should not be a formula to calculate the area of a circle with a diameter of 68 miles. What is the formula for the area of a circle, given its radius? \(Area = \pi r^{2}\) What is the formula for the area of a circle, given its diameter? \(Area = \pi \left(\dfrac{d}{2}\right)^{2} = \dfrac{\pi}{4}d^{2}\) What is the formula for the area of a circle, given its Circumference? \(Area = \pi \left(\dfrac{\dfrac{C}{\pi}}{2}\right)^{2} = \pi\left(\dfrac{C}{2\pi}\right)^{2} = \dfrac{C^{2}}{4\pi}\) Now, if you give me a specific diameter or radius, we are in business and we can just substitute to provided values. This is what a formula does. If you have a "formula" to solves a single problem, something is seriously wrong. What is the formula for the area of a circle, given its area?!